OpenStudy (anonymous):

I have an equation like this: 2pi(r)^2+14pi(r)-231=0 How can I find the value of r?

OpenStudy (anonymous):

use pi =22/7 and then it will com in in simpler terms...

OpenStudy (anonymous):

further solve the quadratic...

OpenStudy (anonymous):

Could you please show me?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

(44/7)(r^2)=A (308/7)(r)=B -231=C for the quadratic formula, let x=r

OpenStudy (anonymous):

Couldn't I just let x = pi x r ??

OpenStudy (anonymous):

@ganeshie8 I tried to do that but it gave me a different answer

OpenStudy (anonymous):

Ax^2+Bx+C=0 That't the standard form. the x, in this case is r. The variable changed, but that shouldn't affect the value.

OpenStudy (anonymous):

Are you typing it in on your calculator incorrectly?

OpenStudy (anonymous):

I'm doing it manually

ganeshie8 (ganeshie8):

wolfram has the correct quadratic formula format look it up

OpenStudy (anonymous):

So I can't let x = pi x r ??

OpenStudy (anonymous):

No

ganeshie8 (ganeshie8):

\( \color{red}{2\pi} r^2+\color{red}{14\pi} r \color{red}{-231}=0\)

ganeshie8 (ganeshie8):

why do you want to let x = pi x r ?

OpenStudy (anonymous):

So that it takes the quadratic form of ax^2 + bx + c

ganeshie8 (ganeshie8):

its already in the quadratic form : \(\color{red}{2\pi} r^2+\color{red}{14\pi} r \color{red}{-231}=0 \) \(a = \color{red}{2 \pi}\) \(b = \color{red}{14 \pi}\) \(c = \color{red}{-231}\)

ganeshie8 (ganeshie8):

And yes, \(\color{red}{\pi}\) is just a number

ganeshie8 (ganeshie8):

if u prefer, u may put \(x = r\)

OpenStudy (anonymous):

Thank you for the explanation :)

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

I have one more question Where did 44/7 r^2 come from?

OpenStudy (anonymous):

I just multiplied it by 2.

OpenStudy (anonymous):

Could you please show me?

OpenStudy (anonymous):

2(22/7) 2*22=44 44/7

OpenStudy (anonymous):

There ya go.

OpenStudy (anonymous):

And the 22/7?

OpenStudy (anonymous):

That's what pi represents as a fraction.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

No problem! Anytime! :)